OpenStudy (anonymous):

Tony and Belinda have a combined age of 56. Belinda is 8 more than twice Tony’s age. How old is each?

OpenStudy (anonymous):

x+2x+8=56 3x+8=56

OpenStudy (shubhamsrg):

let ages be t and b,,of tony and belinda respectively so you're given t + b=56 and 2t + 8 = b just solve for t and b..

OpenStudy (anonymous):

x=Tonys age

OpenStudy (anonymous):

Lets say T=tony`s age and B=belinda`s age So you have T+B = 56 and B = 2T+8 From the first equation: B = 56-T So 56-T = 2T+8 So 48 - T = 2T So 48 = 3T So T = 16 And so if T=16 then B= 56-16 = 40

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

your welcome!